OpenStudy (anonymous):

Factor: d2 – 144 and please explain how you did it if that isn't too much?

OpenStudy (whpalmer4):

Both d^2 and 144 are squares, so this is a difference of two squares. \[(x-a)(x+a) = x^2 + ax - ax - a^2 = x^2-a^2\] All you need to do is figure out what the square root of 144 is, and what the square root of d is, and then you have your factors. \[(\sqrt{d^2} - \sqrt{144})(\sqrt{d^2} + \sqrt{144})\]

OpenStudy (whpalmer4):

Sorry, "what the square root of \(d^2\) is"

OpenStudy (whpalmer4):

Learn to recognize this pattern, you'll use it often when factoring.

OpenStudy (anonymous):

So.. (d-12)(d+12)?

OpenStudy (whpalmer4):

Correctamundo!

OpenStudy (whpalmer4):

\[(d-12)(d+12) = d^2 -12d + 12d -144 = d^2 - 144\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (whpalmer4):

You're welcome. Like I said, keep your eyes open for other opportunities to use that trick!

Directrix (directrix):

Difference of Two Squares Study Aid